Ques) A coil of current crrying Nichrome wire is immersed in aliquid. When the potetial difference across the coil is 12 V and the current through the coil is 5.2 A., the liquidm evaporated at the steady rate of 21mg/sec. Calculate the heat of  Vapourization of the liquid.

Grade:12

3 Answers

AskIITians Expert PRASAD IIT Kharagpur
14 years ago

Power consumed = V*I = 12 *5.2 = 62.4 J/s.

Rate of evapoation = 21 mg/s = 2.1 * 10-5  kg /s

Power consumed = Heat gaines by water per second

Heat of Vaporization, H  = 62.4 / (2.1 * 10-5)  =  2971 KJ/kg =710 Kcal / kg.

Ramesh V
14 years ago

heat generated in coil (H) = I2Rt

given V =12 and I = 5.2 A , by ohms law, R = V/I = 12/5.2 =2.3 ohms

dH/dt = I2R  = 62.2 J/sec

let specific heat of vapourisation is:  L J/g

rate of evaporation is : m' = 21x10-3 g/sec

so, dH/dt = m' * L 

L = (dH/dt)/ m'

L= 62.2 / 21x10-3

L = 2961.5 J/gram

----

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch

Vanya Saxena
14 years ago

Heat generated per sec due to the current carrying coil,

                                                                                        H=VI  =>H =12*5.2 J/sec

 This heat will be equal to the latent heat of vapourization as the same is responsible for the vapourization of the liquid,

                                                       VI=mL      =>12*5.2=21/1000*L

                                                                 L=29.71*10^2 J/Kg